# How Does A Text Preprocessing Pipeline Affect Ontology Syntactic Matching?

### v1

@@ I Introduction
While a number of OM systems use the text preprocessing pipeline for syntactic OM, few studies explain why a specific method is selected for a certain OM task. Our study tackles the challenge in two ways. Firstly, we conduct a comprehensive experimental analysis of the text preprocessing pipeline in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline in OM tasks at syntactic levels. In each phase, a text preprocessing method is evaluated for its correctness and completeness. Secondly, we propose a novel context-based pipeline repair approach for syntactic OM. The method offers a customised way to fine-tune the text preprocessing pipeline for each domain-specific OM task and shows promising results for repairing false mappings. We also discuss the non-deterministic nature of using LLMs for text preprocessing and highlight the importance of utilising the classical text preprocessing pipeline. Specifically, this paper makes the following contributions:
@@ III Preliminaries: OM task formulation
IV Analysis of Text Preprocessing Pipeline
In this section, we illustrate the procedure used to analyse the text preprocessing pipeline in ontology syntactic matching and our evaluation results of the effect of different text preprocessing in the OAEI repositories.
@@ V Context-based Pipeline Repair
Experimental results in the last section demonstrate that only Phase 1 text processing (i.e. Tokenisation and Normalisation) helps with syntactic OM. The use of Phase 2 text preprocessing (i.e. Stop Words Removal and Stemming/Lemmatisation) does not improve performance and may even have negative impacts. In this section, we explore the underlying cause of this phenomenon. Furthermore, we introduce an intuitive context-based pipeline repair approach and demonstrate its effectiveness on the same datasets we analysed.
V-A Motivation
@@ VI Text Preprocessing in the era of LLMs
With the recent advances in LLMs, it becomes questionable whether we still need the classic text processing pipeline with the availability of powerful LLMs. In this section, we demonstrate the non-deterministic nature of LLMs in OM text preprocessing and showcase the potential use of LLMs in OM text preprocessing pipeline repair.
@@ IX Conclusion
The source code for this study is available from https://github.com/qzc438/ontology-nlp. The OAEI track repositories used in this study can be downloaded from https://dwslab.github.io/melt/track-repository (date accessed: 2024-06-01). The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.

### v2

@@ I Introduction
While a number of OM systems use the text preprocessing pipeline for syntactic OM, few studies explain why a specific method is selected for a certain OM task. Our study tackles the challenge in two ways. Firstly, we conduct a comprehensive experimental analysis of the text preprocessing pipeline in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline in OM tasks at syntactic levels. In each phase, a text preprocessing method is evaluated for its correctness and completeness. Secondly, we propose a novel context-based pipeline repair approach for syntactic OM. The method offers a customised way to fine-tune the text preprocessing pipeline for each domain-specific OM task and shows promising results for repairing false mappings. Thirdly, we discuss the non-deterministic nature of LLMs for text preprocessing and highlight the importance of retaining the classical text preprocessing pipeline. Specifically, this paper makes the following contributions:
@@ III Preliminaries: OM task formulation
IV Text Preprocessing Pipeline: Analysis
In this section, we discuss the role of the text preprocessing pipeline in ontology syntactic matching and report our experiments studying the effect of different text preprocessing options over the OAEI repositories.
@@ V Context-based Pipeline Repair
Experimental results in Section IV demonstrate that only Phase 1 text processing (i.e. Tokenisation and Normalisation) helps with syntactic OM. The use of Phase 2 text preprocessing (i.e. Stop Words Removal and Stemming/Lemmatisation) does not improve performance and may even have negative impacts. In this section, we explore the underlying cause of this behaviour. Furthermore, we introduce an intuitive context-based pipeline repair approach and demonstrate its effectiveness on the same datasets.
V-A Causal analysis
@@ VI Text Preprocessing in the era of LLMs
With the recent advances in LLMs, we ask whether we still need the classic text processing pipeline. In this section, we demonstrate the non-deterministic nature of LLMs in OM text preprocessing and showcase the benefits of LLMs in a repair function inside the OM text preprocessing pipeline.
@@ IX Conclusion
The source code for this study is available from https://github.com/qzc438/ontology-nlp. The OAEI track repositories used in this study can be downloaded from OAEI MELT at https://dwslab.github.io/melt/track-repository (date accessed: 2024-06-01). The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.
@@ Acknowledgment
Appendix A Supplementary Material
Appendix B Extended Experiment Details
B-A Analysis of Text Preprocessing Pipeline
Fig. 121314, and 15 show the details of the experiment to compare the text preprocessing pipeline in syntactic OM.
Figure 12: Base without Preprocessing (B) vs. Tokenisation (T).
Figure 13: Tokenisation (T) vs. Normalisation (N).
Figure 14: Normalisation (N) vs. Stop Words Removal (R).
Figure 15: Stop Words Removal (R) vs. Stemming/Lemmatisation (S/L).
  1. (1)
    For Phase 1 text preprocessing methods (Tokenisation and Normalisation), most of the data points located above the equivalent line in Precision, Recall, and F1 Score indicate that they help syntactic OM.
  2. (2)
    For Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation): Some data points are located above the equivalent line in Recall, but the majority of data points located below the equivalent line in Precision and F1 Score indicate that they do not help syntactic OM.
Fig. 16 shows the details of the experiment to compare Stemming (S) and Lemmatisation (L). Most of the data points located above the L=S line in Precision and F1 Score indicate that using Lemmatisation is better than Stemming in syntactic OM (assuming that the post hoc correction is excluded).
Figure 16: Stemming (S) vs. Lemmatisation (L).
Fig. 1718, and 19 show the details of the experiment to compare different stemmers.
Figure 17: Porter Stemmer (SP) vs. Snowball Stemmer (SS).
Figure 18: Porter Stemmer (SP) vs. Lancaster Stemmer (SL).
Figure 19: Snowball Stemmer (SS) vs. Lancaster Stemmer (SL).
  1. (1)
    For Porter Stemmer vs. Snowball Stemmer, all data points located in the equivalent line indicate that there is no difference using Porter Stemmer and Snowball Stemmer in syntactic OM.
  2. (2)
    For Porter/Snowball Stemmer vs. Lancaster Stemmer, most of the data points the data points located above the equivalent line in Precision and F1 Score indicate that Porter/Snowball Stemmer is more effective than Lancaster Stemmer in syntactic OM.
Fig 20 shows the details of the experiment to compare Lemmatisation vs. Lemmatisation + POS Tagging. All data points located in the equivalent line indicate that using POS Tagging in Lemmatisation does not help syntactic OM.
Figure 20: Lemmatisation (L) vs. Lemmatisation + Part-Of-Speech Tagging (LT).
B-B Context-based Pipeline Repair
Fig. 2122232425, and 26 show the details of the experiment to consider the benefit of using context-based pipeline repair in Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation).
  1. (1)
    For Precision, most of the data points are located above the equivalent line, indicating that context-based pipeline repair significantly improves matching correctness.
  2. (2)
    For Recall, most of the data points are located in the equivalent line, and only a few data points are located below the equivalent line, indicating that context-based pipeline repair slightly reduces matching completeness.
  3. (3)
    For F1 Score, most of the data points are located above the equivalent line, indicating that context-based pipeline repair also improves overall matching performance.
  4. (4)
    Experimentally, the wider ellipse around the data points indicates that the matching performance improvement ranking in Phase 2 text preprocessing methods is Stemming (S) > Lemmatisation (L) > Stop Words Removal (R).
  5. (5)
    Experimentally, the wider ellipse around the data points indicates that the matching performance improvement ranking in different stemmers is Lancaster Stemmer (SL) > Porter Stemmer (SP) = Snowball Stemmer (SS).
  6. (6)
    Experimentally, the same ellipse around the data points indicates that the matching performance improvement ranking in lemmatisation is Lemmatisation (L) = Lemmatisation + POS Tagging (LT).
Figure 21: Comparison of using and without using context-based repair in Stop Words Removal (R).
Figure 22: Comparison of using and without using context-based repair in Porter Stemmer (SP).
Figure 23: Comparison of using and without using context-based repair in Snowball Stemmer (SS).
Figure 24: Comparison of using and without using context-based repair in Lancaster Stemmer (SL).
Figure 25: Comparison of using and without using context-based repair in Lemmatisation (L).
Figure 26: Comparison of using and without using context-based repair in Lemmatisation + POS Tagging (LT).